# Jiashu Tao, and Reza Shokri

### v1

@@ 3.1 The Original Robust Membership Inference Attacks (RMIA)
where $\theta$ is the target model, $x$ is the target query, $z$ is population data, and $\gamma\geq 1$ is the threshold. It counts the proportion of ”similar” data the target $x$ dominates. The score is a discrete value, whose precision depends on $|Z|$. More details of RMIA can be found in Appendix A.
@@ 4.1 From Sequences to Tokens
So far, membership inference and privacy risks for LLMs have been defined on the sequence level, i.e., whether a given sequence is a member. The majority of the LLM MIAs aim to compute a score on each sequence [1, 15] based on its perplexity. However, delving into the mechanisms of LLMs, we can quickly realize that a sequence is not one single output, but an ordered list of outputs. For example, given training sequence is $\mathbf{x}=\{x_{1}x_{2}\dots x_{k}\}$, the LLM $\theta$ optimizes the losses $\ell(x_{2},\theta(x_{1})),\ell(x_{3},\theta(x_{1}x_{2})),\dots,\ell(x_{k},\theta(x_{1}\dots x_{k-1}))$. Each sequence is more than one training sample; it resembles a dataset containing k1k-1 training (subsequence, label) pairs. To properly adapt existing MIAs to LLMs, we should treat each token generation step, which ”labels” each ”prefixal” subsequence (subsequences from the start), as one prediction and compute its membership likelihood. In this way, for any sequence of length $k$, the LLM goes through k1k-1 prediction steps, and we should obtain k1k-1 membership scores. In comparison, the existing framework only computes a single membership signal for each sequence, which is a highly compressed signal, losing rich information at each token position.
@@ 6.2 Solving Sequence-Level Membership Inference with Token-Level Membership Signals
We observe that our token-level InfoRMIA, although not specifically designed for sequence-level membership inference, achieves the strongest membership inference performance (Table 3, 4) even when using only a single, less ideal reference model. In Tables 68, we report results obtained by using the step-1 checkpoint of the target model as the reference and observe minimal change in utility. We further show that our method is the strongest reference-based MIA for pretrained LLMs, outperforming prior reference-model approaches (Ref). We also show that our method is the strongest reference-based MIA for pretrained LLMs, compared to Ref. We also find that simple averaging outperforms the min-$k$ aggregation when targeting high true-positive rates at low false-positive rates (TPR at small FPR). This is expected, as min-$k$ aggregation essentially acts as a non-member detector: non-members tend to contain more low-probability tokens [18].Consequently, min-$k$ is less suitable for high precision member detection with minimal error rates. However, when evaluating using AUC (Table 5), the ordering reverses. This aligns with the argument by Carlini et al. [2] that AUC can be misleading as a privacy metric. Nonetheless, many recent works evaluating on MIMIR still report only AUC in their main text, which may encourage a suboptimal trend for future development of LLM MIAs.
@@ Appendix A The Original RMIA
In simple terms, RMIA counts the proportion of ”similar” data the target $x$ dominates. In practice, the test statistic is written as
@@
Note that the formulation in Eqn 10 makes the membership score a discrete value, whose granularity depends on the size of $Z$. Intuitively, the more $z$ data RMIA uses, the finer the ”bins” become, and the more distinguishing and precise the signal gets. Empirically, Zarifzadeh et al. [26] have also reported this relationship between the size of $Z$ and the attack performance. The empirical insight was that using $Z$ of about $10\%$ of the training set size is sufficient. However, for LLMs, even $10\%$ represents an astronomical number of samples.
@@ D.1 AG News
The ai4privacy dataset we used is the pii-masking-300k variant, that can be access at https://huggingface.co/datasets/ai4privacy/pii-masking-300k. It is divided into two parts: OpenPII-220k and FinPII-80k. The FinPII has additional classes that are specific to the Finance and Insurance domains. In this dataset, there is a ”privacy_mask” column that marks the beginning and end location for each piece of private information. Thus, we can use this information to categorize each token as private or non-private. It also assigns a type to private substrings, such as last names or email addresses, enabling us to do more interesting analysis.
@@ D.3.2 Why ReCaLL was Excluded in Our Table
The ReCaLL attack pushed to the MIMIR benchmark is a simplied version (according to the author), and very problematic. There is an implicit information leakage about the membership labels in crafting the attack signals, making the result unreliable and unfair. By right, the attacker should not be able to distinguish any non-member from any member, which means the attacker has no information that can be used to tell non-members apart from members before the attack. However, the ReCaLL attack in MIMIR explicitly uses non-members in the evaluation set as its ”non-member” prefix to be prepended to all sequences, making the attacker aware of the membership label of certain non-members. Although the label information is not explicitly used in the attack, it is an implicit information leak that can be used to tell apart the two sets. Hence, the current implementation of the ReCaLL attack in MIMIR is not correct. This is why we did not include it in our tables.

### v2

@@ 3.1 The Original Robust Membership Inference Attacks (RMIA)
where $\theta$ is the target model, $x$ is the target query, $z$ is population data, and $\gamma\geq 1$ is the threshold. It counts the proportion of “similar” data the target $x$ dominates. The score is a discrete value, whose precision depends on $|Z|$. More details of RMIA can be found in Appendix A.
@@ 4.1 From Sequences to Tokens
So far, membership inference and privacy risks for LLMs have been defined on the sequence level, i.e., whether a given sequence is a member. The majority of the LLM MIAs aim to compute a score on each sequence [1, 15] based on its perplexity. However, delving into the mechanisms of LLMs, we can quickly realize that a sequence is not one single output, but an ordered list of outputs. For example, given training sequence is $\mathbf{x}=\{x_{1}x_{2}\dots x_{k}\}$, the LLM $\theta$ optimizes the losses $\ell(x_{2},\theta(x_{1})),\ell(x_{3},\theta(x_{1}x_{2})),\dots,\ell(x_{k},\theta(x_{1}\dots x_{k-1}))$. Each sequence is more than one training sample; it resembles a dataset containing k1k-1 training (subsequence, label) pairs. To properly adapt existing MIAs to LLMs, we should treat each token generation step, which “labels” each “prefixal” subsequence (subsequences from the start), as one prediction and compute its membership likelihood. In this way, for any sequence of length $k$, the LLM goes through k1k-1 prediction steps, and we should obtain k1k-1 membership scores. In comparison, the existing framework only computes a single membership signal for each sequence, which is a highly compressed signal, losing rich information at each token position.
@@ 6.2 Solving Sequence-Level Membership Inference with Token-Level Membership Signals
We observe that our token-level InfoRMIA, although not specifically designed for sequence-level membership inference, achieves the strongest membership inference performance (Table 34) even when using only a single, less ideal reference model. In Tables 68, we report results obtained by using the step-1 checkpoint of the target model as the reference and observe minimal change in utility. We further show that our method is the strongest reference-based MIA for pretrained LLMs, outperforming prior reference-model approaches (Ref). We also show that our method is the strongest reference-based MIA for pretrained LLMs, compared to Ref. We also find that simple averaging outperforms the min-$k$ aggregation when targeting high true-positive rates at low false-positive rates (TPR at small FPR). This is expected, as min-$k$ aggregation essentially acts as a non-member detector: non-members tend to contain more low-probability tokens [18].Consequently, min-$k$ is less suitable for high precision member detection with minimal error rates. However, when evaluating using AUC (Table 5), the ordering reverses. This aligns with the argument by Carlini et al. [2] that AUC can be misleading as a privacy metric. Nonetheless, many recent works evaluating on MIMIR still report only AUC in their main text, which may encourage a suboptimal trend for future development of LLM MIAs.
@@ Appendix A The Original RMIA
In simple terms, RMIA counts the proportion of “similar” data the target $x$ dominates. In practice, the test statistic is written as
@@
Note that the formulation in Eqn 10 makes the membership score a discrete value, whose granularity depends on the size of $Z$. Intuitively, the more $z$ data RMIA uses, the finer the “bins” become, and the more distinguishing and precise the signal gets. Empirically, Zarifzadeh et al. [26] have also reported this relationship between the size of $Z$ and the attack performance. The empirical insight was that using $Z$ of about $10\%$ of the training set size is sufficient. However, for LLMs, even $10\%$ represents an astronomical number of samples.
@@ D.1 AG News
The ai4privacy dataset we used is the pii-masking-300k variant, that can be access at https://huggingface.co/datasets/ai4privacy/pii-masking-300k. It is divided into two parts: OpenPII-220k and FinPII-80k. The FinPII has additional classes that are specific to the Finance and Insurance domains. In this dataset, there is a “privacy_mask” column that marks the beginning and end location for each piece of private information. Thus, we can use this information to categorize each token as private or non-private. It also assigns a type to private substrings, such as last names or email addresses, enabling us to do more interesting analysis.
@@ D.3.2 Why ReCaLL was Excluded in Our Table
The ReCaLL attack pushed to the MIMIR benchmark is a simplied version (according to the author), and very problematic. There is an implicit information leakage about the membership labels in crafting the attack signals, making the result unreliable and unfair. By right, the attacker should not be able to distinguish any non-member from any member, which means the attacker has no information that can be used to tell non-members apart from members before the attack. However, the ReCaLL attack in MIMIR explicitly uses non-members in the evaluation set as its “non-member” prefix to be prepended to all sequences, making the attacker aware of the membership label of certain non-members. Although the label information is not explicitly used in the attack, it is an implicit information leak that can be used to tell apart the two sets. Hence, the current implementation of the ReCaLL attack in MIMIR is not correct. We will include it once the official implementation is fixed.